40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Oct 14 23:33:45 2010, Suresh, Update, IOO, 2W NPRO laser output power versus temperature tempscan.pngScan2010.zip
    Reply  Fri Oct 15 00:15:52 2010, Koji, Update, IOO, 2W NPRO laser output power versus temperature 
Message ID: 3725     Entry time: Thu Oct 14 23:33:45 2010     Reply to this: 3726
Author: Suresh 
Type: Update 
Category: IOO 
Subject: 2W NPRO laser output power versus temperature 

Steve measured an apparent power drop in the 2W NPRO output from 2.1W to 1.6W(elog entry no 3698) at 2.1A of diode current in the laser (elog entry:  2822).  It was later noticed that the laser temperature was set to about 45 degC while the initial calibration was done at 25 deg C.  

It was felt that the recent power drop may have something to do with the increase in the operating temperature of the laser from 25 to 45 deg C.  Therefore the laser was returned to 25 deg C and the power output was remeasured and found to be 2.1W as it was at the begining(elog entry:3709)

It was also noticed that returning the laser to 25 deg. C resulted in a loss of efficiency in coupling to the PMC.  We suspected that this might be due to multimode operating conditions in the laser at particular operating temperatures.  In order to see if this is indeed the case the laser power output was observed as a function of temperature.  We do notice a characteristic saw-tooth shape which might indicate multimode operation between 39 and 43 deg C.  It is best to verify this by observing the power fluctuations in the transmitted beam of the stabilised reference cavity.

 

tempscan.png

 

The measurement was made by attenuating the roughly 2W laser beam by a stack of two Neutral Density filfers and then measuring the transmitted light with the PDA36A photodetector.  This was because both the power meters used in the past were found to have linear drifts in excess of 30% and fluctuations at the 10% level. 

 

 

Attachment 2: Scan2010.zip  2 kB  Uploaded Fri Oct 15 01:10:50 2010
ELOG V3.1.3-